40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 28 03:01:24 2015, Jenne, Update, LSC, Transitioned DARM to AS55Q 
    Reply  Wed Jan 28 14:19:02 2015, rana, Update, LSC, Transitioned DARM to AS55Q 
Message ID: 10947     Entry time: Wed Jan 28 03:01:24 2015     Reply to this: 10949
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Transitioned DARM to AS55Q 

[Jenne, Diego]

Tonight we were able to transition DARM from DC transmission signals to AS55Q/(TRX+TRY).  That's about as far as we've gotten though.

Here are the details:

  • Set the ASDC->MICH matrix element such that the MICH fringes were 0-1.  For some reason this number seems to change by ~10% or so each night.
  • Followed main carm_cm_up script, although stopped at MICH offset of 25% (mostly because I forgot to let it go to 50% - no fundamental reason)
  • So, MICH was at 25% (with a + for the gain accidentally, even though we decided yesterday that - was better), arm powers were about 1.1 or so.
  • Took transfer functions driving DARM and looking at normalized AS55Q, and driving CARM looking at normalized REFL11I.
    • There is still not a lot of coherence in the CARM->REFL11I case, so we decided to stick with DARM for starters.
    • The TF between DARM and AS55Q looked really nice!
  • Prepared the unused DARM error signal row, including an offset before the blend matrix.
  • Transitioned over to normalized AS55Q.
    • This left the DARM servo filterbank with a zero digital offset.
    • But, the error signal had an offset before it got to the DARM filter bank.
      • This offset does not have any ramping (I don't know how to do that when building a model), so it's not as nice for reducing an offset.
      • Maybe we can, after transitioning to the new signal, move the offset to the DARM servo filterbank?
  • Was reducing the DARM offset so that we were at the true AS55Q zero crossing.
  • Saw that the UGF servo lines were starting to get a bit lost in the noise, so I increased the DARM's amplitude.
    • I don't know if the UGF servo was already too far gone and increasing the SNR couldn't recover it, or if I was driving too hard and directly kicked myself out of lock.  Either way, we lost lock.

The carm_cm_up script now should get all the way to this point by itself, although occasionally the PRMI part will lose lock (but not the arms), so you have to go back to the 3nm CARM offset and relock the central part.  I have written a little "relockPRMI.sh" script that lives in ..../scripts/PRFPMI/ that will take care of this for you.  

We were able to transition DARM to AS55Q a total of 3 or so times tonight.  The first time was with the + MICH gain, and the rest of the times were with - MICH gain.  The carm_up script now asks for a sign for the MICH gain just after asking for a CARM offset sign. 

I think that we understand all of our locklosses from these states.  Twice (including the time described above) the UGF lines got lost in the noise, and the UGF servos went crazy.  Once the PRCL loop rang up, because a filter that wasn't supposed to be on was on.  This was a terrible filter that I had made a long time ago, and was never supposed to be part of the locking sequence, but it was getting turned on by a restore script, and kept eating our phase.  Anyhow, I have deleted this terrible boost filter so it won't happen again (it was called "boost test", which may give you an idea of how non-confident I was in its readiness for prime-time).  The last time of the night I must not have been quite close enough in CARM offset, so we should probably check with a TF before making this last jump.

Diego wrote a nifty burt restoring script that will clear out all the elements of the input matrix and the normalization matrix for a row that you tell it (i.e. DARM_A will clear out all the elements in the first row of those 2 matrices).  This is useful for the switches back and forth between the _A and _B signals, to make sure that everything is in order.  So, I now run those clear scripts, then put in the elements that I want for the next step.  For example, DARM initially locks with ALS using the A row.  Then, DARM transitions to the B row for DC transmission.  Then, I prepare the A row for AS55Q locking, and I don't want any elements accidentally left from the ALS signal.  It lives in ..../scripts/LSC/InputMatrix/

 

Thoughts for tomorrow:

Daytime re-commission the Xarm ASS.

Nighttime try to get back to DARM on AS55Q and push farther forward. 

 

ELOG V3.1.3-